OpenStudy (anonymous):

Solve 9 - (3 - x) = 14

OpenStudy (btaylor):

9 - 3 +x = 14 6 + x = 14 x = 8

OpenStudy (btaylor):

@zac629 you forgot to distribute the -1 to the three. If you checked your answer, it wouldn't be right.

OpenStudy (anonymous):

9 - 14 = 3 - x -5 = 3 - x x = 3 +5 x = 8 is the answer..

OpenStudy (anonymous):

o sorry haha